A rectangular swimming pool 30 feet by 15 feet is surrounded by a uniform concrete sidewalk 5 feet wide. The pool is positioned inside a fenced- in circular grassy region of diameter 60 feet. a. Find the surface area of the water. b. Find the area of the concrete sidewalk. c. Find the area of the grass.
Answer by mananth(16946)

surface area of water = 30*15 = 450 sq.ft
Surface area including side walk = (30+10)*(15+10)= 1000 sq.ft
So Area of sidewalk = 1000-450= 550 sq.ft
...
Fencing Area = pi * r^2
Radius = 30 ft.
Area = 3.14*30^2= 2826 sq.ft
Area of grassy region = 2826-1000 = 1826 sq.ft
Draw a sketch and the problem will be easy to solve